Nova Company’s total overhead cost at various levels of activity are presented below:
Month Machine-Hours Total Overhead Cost
     April 52,000           $192,320        
     May 42,000           $166,720        
     June 62,000           $217,920        
     July 72,000           $243,520        

Assume that the total overhead cost above consists of utilities, supervisory salaries, and maintenance. The breakdown of these costs at the 42,000 machine-hour level of activity is:

  Utilities (variable) $ 58,800  
  Supervisory salaries (fixed) 43,000
  Maintenance (mixed) 64,920
  Total overhead cost $ 166,720

Nova Company’s management wants to break down the maintenance cost into its variable and fixed cost elements.

Express the company’s total overhead cost in the linear equation form Y = a + bX. (Do not round your intermediate calculations. Round the "Variable cost per unit" to 2 decimal places.)

     

What total overhead cost would you expect to be incurred at an activity level of 47,000 machine-hours? (Do not round intermediate calculations.)
Answer #1

1)Equation :   y = 2.56 x +59200

where y =total cost ,x = number of machine hours

Variable cost per MH Fixed cost
utilities 58800/42000=1.4
supervisory salaries 43000
maintenance 1.16 16200
Total cost 2.56 59200

**maintenance cost:

At highest activity level of 72000 MH =243520-(72000*1.4)utilities cost - 43000 = 99720

At lowest activity level of 42000 MH =64920

variable cost per MH = Change in cost /change in MH

                =[99720-64920]/[72000-42000]

                = 34800/30000

               = $ 1.16 MH

FIxed cost at highest activity = 99720 - [72000*1.16]

            = 16200

2)Total overhead cost at 47000MH = [2.56*47000]+59200

                     = 120320+59200

                          = 179520